OpenStudy (anonymous):

Convert the equation from Rectangular to polar form: x^2 - y^2 = 9

OpenStudy (jdoe0001):

same as the one you did the other day, \(\bf x=rcos(\theta)\qquad y=rsin(\theta)\)

OpenStudy (jdoe0001):

then take common factor on the left-side and solve for "r"

OpenStudy (anonymous):

Right, thank you. So it would be \[r^2 = 9 / (\cos^2 (\theta) - \sin^2 (\theta)\]

OpenStudy (jdoe0001):

well... that's correct, but solving for "r" will give you \(\bf r=\sqrt{\cfrac{9}{cos^2(\theta)-sin^2(\theta)}}\)